Pedro is five less than three times as old as his sister Sofia. John is twice as old as Pedro. Together their ages add up to 45. Which of these could represent Pedro's, Sofia's, or John's age.

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)


Pedro is five less than three times as old as his sister Sofia. John is twice as old as Pedro. Together their ages add up to 45. Which of these could represent Pedro's, Sofia's, or John's age.
Pedro is five less than three times as old as his sister Sofia
p= 3s-5
John is twice as old as Pedro.
j=2p
therefore j= 2(3s-5)= 6s-10
Together their ages add up to 45.
j+p+s=45
substitute j & p
6s-10+3s-5 +s=45
10s = 60
s=6
p= 13
j=2p
j=26

Answer by josgarithmetic(39618)

Pedro, Sofia, and John ---
PEDRO             3x-5

SOFIA               x

JOHN             2(3x-5)

TOTAL_SUM           45